OpenStudy (anonymous):

A cone has a radius of 8 inches and a height of 12 inches. What is the exact volume of the cone? A. 32π in3 B. 96π in3 C. 256π in3 D. 768π in3

OpenStudy (acxbox22):

\[V=\frac{ 1 }{ 3 }\pi r^2h\]

OpenStudy (acxbox22):

v=1/3(8)^2pi(12) v=4(64pi) v=??

OpenStudy (anonymous):

c?

OpenStudy (acxbox22):

yes,good job

OpenStudy (anonymous):

Thank you ! :)